OpenStudy (moonlitfate):

The populations P (in thousands) of a certain city from 2000 through 2007 can be modeled by P = 343.8e^(k*t), where t represents the year, with t = 0 corresponding to 2000. In 2004, the population of the city was about 392,000. (a) Find the value of k. (Round your answer to five decimal places.)

OpenStudy (moonlitfate):

I tried to solve for k and I come up with k = 1.75974. Looks like I did something wrong, though. Can someone help me find out where that was?

OpenStudy (anonymous):

hi moonlit. question: what did you do to solve for k?

OpenStudy (moonlitfate):

Gimme a second and I'll type it all out? ^^

OpenStudy (anonymous):

k

OpenStudy (moonlitfate):

\[392000 = 343.8e ^{k4} \] \[\frac{ 392000 }{ 343.8 } = e ^{4k}\] \[1140.1978 = e ^{4k}\] \[\ln 1140.1978 = 4k\] \[\frac{\ln 1140.1978}{ 4 } = k\] \[k \approx 1.75974\]

OpenStudy (anonymous):

your process is absolutely right. :)

OpenStudy (moonlitfate):

For some reason, it's not correct. It doesn't make sense at all. o:

OpenStudy (anonymous):

what is the given answer?

OpenStudy (moonlitfate):

I haven't been given the answer, it's just marked wrong. And only have one more shot at it.

OpenStudy (anonymous):

are you sure you copied it correctly?

OpenStudy (moonlitfate):

Problem is in the attachment

OpenStudy (anonymous):

i would talk to your teacher because the answer is correct.

OpenStudy (anonymous):

incidentally, i am a california credentialed math teacher. :)

OpenStudy (anonymous):

i could be wrong. it has happened before. but your process and numbers match mine exactly.

OpenStudy (anonymous):

their sig figs could be different. I have found that computer generated problems can sometimes be wrong.

OpenStudy (moonlitfate):

Hm, I'll try e-mailing my professor and taking a screen capture of what I did to solve for k. It's just odd.

OpenStudy (anonymous):

i would do that first. ask if the process is correct?

OpenStudy (anonymous):

it is odd. try your professor. see if you can get more info. :) your solution is correct. looked at it 3 times now.